In a car race , car A takes time t less than car B and passes the finishing point with a velocity v more than the velocity with which car B passes the point. Assuming that the cars start fromrest and travel with constant accelerations a1 and a2 show that            v = (a1a2)1/2

Grade:10

1 Answers

Tarsh
5 years ago
Let s be the distance travelled by each carvA=2a1s−−−−√vA=2a1sAlso vB=2a2s−−−−√vB=2a2s t1=2sa1−−−√t1=2sa1t2=2sa2−−−√t2=2sa22a1s−−−−√−2a2s−−−−√=v02a1s−2a2s=v0and2sa2−−−√−2sa1−−−√=t02sa2−2sa1=t0v0t0=a1−−√−a2−−√1a2√−1a1√v0t0=a1−a21a2−1a1=a1a2−−−−√